OpenStudy (anonymous):

Simplify: 2√3-4√2+6√3+8√2 a) 8√3+4√2 b) 12√5 c) -4√3-6√2 d) 8√6+8

OpenStudy (igreen):

You have to be more specific..what numbers are in the square roots? Is it: \(\sqrt{3-4} \sqrt{2+6} \sqrt{3 + 8} \sqrt2\) or is it\(\sqrt{3} - 4 \sqrt2 + 6 \sqrt3 + 8 \sqrt2\)?

OpenStudy (anonymous):

I don't know, this is all they show me. I have no idea how to do this or I would'nt ask.

OpenStudy (igreen):

Can you take a printscreen of the question?

OpenStudy (anonymous):

I typed it here exactly the way it looks.

OpenStudy (igreen):

Then I have no idea how to solve it.

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):